find the roots of the equation ((3z-1)^4)+((z-2)^4)=0 where z=x+iy

Grade:11

1 Answers

prateek
7 years ago
(3z-1)^4= -(Z-2)^4
3z-1=2-Z
4z=3
Z=3/4
NOW U CAN SOLVE THIS 
PLEASE CHECK OUT COURSES WE PROVIDE
 I HOPE YOU WILL FOUND THISANSWER HELPFUL THANKS FOR POSTING YOUR QUERIES ON ASKIITIANS.COM